# DART-Eval: A Comprehensive DNA Language Model Evaluation Benchmark on Regulatory DNA

### v1

@@ E.5 Predicting counterfactual effects of regulatory genetic variants
Probed Caduceus -0.0044 0.0040 0.5124 0.0848 0.0003
DNABERT-2 0.0064 0.0111 0.5024 0.0810 0.2536
GENA-LM -0.0067 0.0082 0.5149 0.0836 2.2893e-5
HyenaDNA 0.0121 0.0139 0.5658 0.0937 4.0326e-73
Mistral-DNA 0.0185 0.0142 0.5018 0.0821 0.2706
Nucleotide Transformer 0.0058 0.0052 0.5248 0.0870 5.4018e-12
Fine-Tuned Caduceus 0.2591 0.2818 0.6498 0.1791 0.000
DNABERT-2 0.1840 0.2185 0.6155 0.1380 1.1660e-220
GENA-LM 0.2011 0.2161 0.6038 0.1285 1.2926e-178
HyenaDNA 0.2653 0.2871 0.6108 0.1233 4.358e-203
Mistral-DNA 0.0849 0.0858 0.5101 0.0841 0.0027
Nucleotide Transformer 0.2299 0.2435 0.6231 0.1542 2.8400e-250
Ab initio ChromBPNet 0.6712 0.6995 0.7716 0.3972 0.000
@@
DNABERT-2 0.0240 -0.0049 0.4756 0.0194 0.9760
GENA-LM 0.0589 0.0292 0.4655 0.0191 0.9975
HyenaDNA -0.0417 -0.0771 0.4672 0.0187 0.9961
@@
DNABERT-2 0.4727 0.5007 0.6307 0.0416 1.390e-26
GENA-LM 0.4141 0.4541 0.6280 0.0396 1.524e-25
HyenaDNA 0.5031 0.5011 0.5729 0.0289 1.6439e-09

### v2

@@ E.5 Predicting counterfactual effects of regulatory genetic variants
Probed Caduceus -0.0209 -0.0084 0.5124 0.0848 0.0003
DNABERT-2 0.0169 0.0106 0.5024 0.0810 0.2562
GENA-LM -0.0087 0.0009 0.5149 0.0836 2.263e-05
HyenaDNA 0.0304 0.0262 0.5658 0.0937 3.920e-73
Mistral-DNA 0.0122 0.0068 0.5018 0.0821 0.2706
Nucleotide Transformer 0.0203 0.0267 0.5248 0.0870 5.336e-12
Fine-Tuned Caduceus 0.4707 0.4945 0.6498 0.1791 0.000
DNABERT-2 0.3899 0.4267 0.6385 0.1524 0.000
GENA-LM 0.3666 0.3772 0.6147 0.1384 1.5595e-217
HyenaDNA 0.3892 0.3958 0.5845 0.1060 3.9080e-119
Mistral-DNA 0.1454 0.1470 0.5101 0.0841 0.0027
Nucleotide Transformer 0.4592 0.4578 0.6225 0.1527 4.5174e-248
Ab initio ChromBPNet 0.6739 0.7033 0.7716 0.3972 0.000
@@
DNABERT-2 0.0239 -0.0044 0.4758 0.0194 0.9756
GENA-LM 0.0589 0.0293 0.4655 0.0191 0.9975
HyenaDNA -0.0416 -0.0771 0.4672 0.0187 0.9961
@@
DNABERT-2 0.5002 0.5303 0.6491 0.0556 5.239e-34
GENA-LM 0.4297 0.4294 0.6409 0.0543 1.450e-30
HyenaDNA 0.3576 0.3304 0.4992 0.0221 0.5272